OpenStudy (anonymous):

x over 8 -9=-2 how do i simplify and solve this problem

OpenStudy (anonymous):

\[\frac{ x }{ 8-9 } = -2\] ?

OpenStudy (anonymous):

x/(8-9)=-2 x=-2(8-9) x=2